# Exploring Radar Data Representations in Autonomous Driving: A Comprehensive Review

### v2

@@ I Introduction
As depicted in Fig. 1, radar sensors have extensive applications in autonomous driving, serving various purposes in the perception of the surrounding environment. Radars mounted on the roof, front and angular directions of an intelligent vehicle are utilized to detect objects in the path and surroundings, providing essential information for collision avoidance and adaptive cruise control [7]. The radar inside the cabin is leveraged to monitor occupant vital signs, left-behind children and driver behaviors (e.g., drowsiness, dangerous maneuvers, health emergencies) in a privacy-protected manner, allowing the autonomous driving system to make alerts as well as safety measures [8, 9, 10, 11]. Furthermore, radar systems deployed alongside roadways can measure the speed of passing vehicles and estimate traffic flow. These traffic data are valuable for traffic optimization, congestion management and intelligent transportation systems, leading to improved road safety and smoother traffic operations [12, 13].
@@ II-A Working Principles
Fig. 3 demonstrates a basic radar system that comprises various analog components, including a synthesizer, a Power Amplifier (PA), transmitters (TX antennas), receivers (RX antennas), a Low Noise Amplifier (LNA), and a mixer, as well as digital components such as an Analog-to-Digital Converter (ADC), a Digital Signal Processor (DSP) and a MicroController Unit (MCU). When the radar starts working, first, the synthesizer generates a linear frequency-modulated pulse called a “chirp”, which is amplified in power and transmitted by the TX antenna. Second, the RX antenna captures the reflected chirp of the target that is amplified with low noise. Third, by combining the RX and TX signals, the mixer produces an Intermediate Frequency (IF) signal, which is then converted into digital values via the ADC. Generally, a radar system contains multiple TX and RX antennas, resulting in multiple IF signals. Information about the target object, such as range, Doppler velocity and azimuth angle, is contained in these IF signals, which can be separated by the DSP using embedding different signal processing algorithms [15]. Based on extracted target information from the DSP and dynamic information from the radar, the MCU serves as a computer to evaluate system requirements and make informed decisions.
@@ II-B1 Parameter Estimation
Range ${R}=\frac{cf_{b}}{2S}$ (1) ${R_{res}}=\frac{c}{2B}$ (2) ${c}$ is the speed of the light ($3*10^{8}m/s$). ${f_{b}}$ is the instantaneous frequency difference at the mixers from TX-chirp and RX-chirp. ${S}$ is the slope of a chirp. ${B}$ refers to the frequency variation range.
Doppler Velocity ${V}=\frac{\lambda\Delta\phi}{4\pi T_{c}}$ (3) ${V_{res}}=\frac{\lambda}{2T_{f}}$ (4) $\lambda$ is the wavelength of the light. ${\Delta\phi}$ is a shift in phase. ${T_{c}}$ is the duration between the chirps. ${T_{f}}$ is cycle time of a frame.
@@ II-B2 RCS Measurement
In terms of mathematical calculations, RCS is a metric that quantifies the ratio between the scattered density in the direction of the radar and the power density intercepted by the object. Since the power is distributed over a sphere, only a small part of this ($4\pi r^{2}$) can be received by the radar. Hence, the expression for RCS takes the form:
@@ III Radar Data Representations
Fig. 6 illustrates the generation process of five different radar data representations. Initially, the raw output from the radar sensor, known as the ADC signal, is hardly interpretable by human observers. Thus, researchers utilize 3D FFTs across the sample, chirp, and antenna dimensions to transform the ADC signal into an image-like representation called the radar tensor. Subsequently, peak detection is employed on the radar tensor to eliminate clutter, resulting in a sparse, point-like representation referred to as the point cloud. By accumulating point clouds over a specific duration and applying grid mapping methods, the grid map representation is generated for the purpose of identifying static objects. Moreover, some researchers perform a Time-Frequency transform following the Range-FFT to extract the micro-Doppler signature, which is utilized for recognizing objects characterized by micro-motion features.
@@
CI4R [66] Micro-Doppler Signature 2020 Classification Radar (77GHz, 24GHz, Xethru) Activities (walking, picking, sitting, crawling, kneeling, limping)
@@ III-A2 Methods
The vital sign-based methods focus on monitoring the physiological signals of occupants inside the vehicle cabin, including heart rate, respiration rate, and presence detection [82, 9, 73, 83, 84]. Signal processing techniques, such as spectrogram analysis, Short Time Fourier Transform (STFT) and Wavelet Transform (WT), are employed to extract the vital sign signals from the radar ADC signal [8, 84]. Machine learning algorithms, such as statistical classifiers or regression models, may be trained using labeled data to estimate vital sign parameters from the extracted features [8].
Radar systems used for cabin vital sign monitoring are generally designed for close-range operation, such as within a vehicle or confined space. They focus on high-resolution detection of small movements and physiological changes, and may use less power and have smaller form factors than road-oriented radar systems. However, vital sign-based methods using machine learning algorithms may have limitations in real-time processing, which is crucial in low-latency applications [85, 86].
@@ III-A3 Discussion
Another limitation is that ADC signals require a developer kit for data acquisition, which makes data processing challenging. Real-time object perception and vital sign monitoring pose computational challenges, requiring efficient hardware and software solutions. Moreover, neural networks processing ADC signals are suitable for classification tasks in closed environments, such as determining the presence or absence of objects and distinguishing between different types of objects [87]. However, in open environments, various distracting factors may lead to inaccurate object feature identification.
@@ III-B2 Methods
2D tensor-based methods typically process 2D input radar data by considering range-Doppler and range-azimuth information. To conduct object classification or detection on radar tensors, image-based network architectures (e.g., Faster R-CNN [90], ResNet [91], U-Net [92]) are applied or modified, as has been demonstrated in various studies using RA tensors [88, 93, 4, 94, 95, 96] or RD tensors [97, 98, 99, 96]. For example, Gao et al[88] initiated the object detection process by defining a fixed-size bounding box for each detected object on the RA tensor. In addition, they extracted the radar data within the bounding box from multiple RA tensors and arranged this data in the form of radar data cubes. With the temporal information and object movement pattern, the data cubes serve as features for classifying different objects, such as pedestrians, cars and cyclists. Dong et al[93] applied ResNet [91] architecture on RA tensors for object detection. They also introduced uncertainty estimation for oriented bounding box localization further to enhance the accuracy of the object detection process. In general, 2D tensor-based methods employ a simplified radar data representation, reducing computational complexity. They can exploit correlations between two dimensions to improve detection and localization performance. However, these methods may only employ part of the available 3D spatial information from the radar. With limited representation, the accuracy of complex scenarios with multiple objects or occlusions may be reduced.
@@
Compared to using 2D tensor, 3D tensor-based and 4D tensor-based methods offer a more comprehensive representation of radar data, leveraging the complete spatial information and allowing for precise object detection, segmentation, and localization. These methods can effectively handle complex scenarios involving multiple targets, partial occlusions, and varying azimuth angles. However, the increased dimensions of the tensor may result in higher computational complexity and system delays.
@@ III-B3 Discussion
Although radar tensors retain more comprehensive information about an object, they also reserve noise and clutter information, which may limit the ability to capture subtle object characteristics or distinguish between similar objects. Furthermore, the radar tensor provides a condensed representation of radar data, reducing the overall data size compared to raw ADC signals. Nevertheless, the radar tensor representation still requires considerable memory storage as well as a large bandwidth, especially for the new 4D radar tensors.
@@ III-C1 Datasets
Conventional 3D radar sensors generate sparse point clouds, as observed in datasets such as nuScenes [52], Zender [41], SeeingThroughFog [54], HawkEye [107], AIODrive [55], RADIal [39], RadarScenes [56], aiMotive [60] and MiliPoint [64]. Recently, 4D radar datasets have emerged, including Astyx [108], VoD [58], TJ4DRadSet [59], WaterScenes [61] and Dual-Radar [63]. While Astyx [108] stands as the first 4D radar point cloud dataset, it is limited by a small data size of only 500 frames. VoD [58] and TJ4DRadSet [59] datasets represent notable advancements in terms of data categories and data size. Meanwhile, these two datasets also incorporate simultaneous LiDAR data, facilitating comparative analysis between 4D radar point clouds and LiDAR point clouds. Interestingly, WaterScenes [61] is a 4D radar dataset focused on objects in waterway environments, which demonstrates the robustness of using 4D radar on water surfaces, particularly in adverse lighting and weather conditions. Dual-Radar [63] incorporates two types of 4D radars, expanding the comparison of different radar performance and further research on practical 4D radar perception algorithms.
@@ III-C2 Methods
Point-based methods [113, 109, 114, 110, 5, 120, 121] process radar point clouds at the individual point level, extracting spatial features directly from these points. These methods generally draw inspiration from LiDAR-based algorithms, such as PointNet [130], PointNet++ [131] and Frustum PointNets [132]. These techniques leverage neural networks with convolutional filters to process radar points directly, considering spatial coordinates as integral features in the network structure. Schumann et al[113] proposed a structure based on PointNet++ for semantic segmentation on radar point clouds. Specifically, the Multi-Scale Grouping (MSG) module in PointNet++ helps group and generate features for a center point and its neighborhoods. They demonstrated that incorporating the RCS value and compensated Doppler velocity significantly improved classification accuracy. Inspired by [113], Danzer et al[109] adopted a two-stage method using PointNets [130, 132] for 2D car detection and segmentation. They considered each point as a proposal and adjusted the proposal size according to the object’s prior knowledge. Then, PointNet and Frustum PointNets are used to classify the proposals and each point in the proposal. Finally, the bounding box prediction is executed only for the proposals that are objects. Point-based methods are good at accurate localization and classification of objects by considering individual radar points. However, they are susceptible to noise and occlusion, which may affect feature extraction accuracy.
@@
Voxel-based methods [133, 111, 134, 135] transform radar point clouds into volumetric representations by discretizing the 3D space into a grid of voxels. Then, 3D convolutions or sparse convolutional neural networks (e.g., VoxelNet [136], SECOND [137]) are leveraged to process the voxels to extract features for perception tasks. For example, GRIF Net [133] initially converts point clouds into voxels. As point clouds typically exhibit sparsity with numerous empty voxels, GRIF Net utilizes the Sparse Block Network (SBNet) [138] to convolve only on masked regions, thereby avoiding unnecessary computations on ineffective blank areas. In LXL [135], the input radar point cloud is initially pillarized as in PointPillars [139]. Then, the pillar representation is fed into SECOND to extract multi-level BEV features with spatial and contextual information. Above all, voxel is an efficient representation for processing dense radar point clouds and is robust to varying point densities and occlusion. However, voxel-based methods lose fine-grained spatial information due to voxelization and discretization. Moreover, the limited resolution of voxel grids may lead to decreased accuracy in object detection, especially for small or distant objects.
@@ III-C3 Discussion
Point cloud representation provides an intuitive 3D spatial structure of the surroundings. Since radar point clouds are generated by filtering techniques, they have advantages in filtering noise and being a lightweight data representation. However, some potential information within the raw data may inevitably be lost [3]. Additionally, some small objects or objects with weak reflections may not be rendered as point clouds.
@@ III-D Grid Map
Leveraging multi-frame point clouds, some researchers transform 3D radar reflections into a BEV pseudo-image called the grid map [113]. There are two primary types of grid maps commonly used in radar-based perception: one is the occupancy-based grid map [140, 141, 142, 143, 144, 145, 146, 147, 148, 149], which represents the obstacles and free-space information; the other is the amplitude-based grid map [140, 141], which displays the reflected power values in that particular area. As shown in Fig. 9(b), an occupancy-based grid map has a clear outline, making it more feasible for outline-based detection tasks (e.g., parking spaces). An amplitude-based grid map, as indicated in Fig. 9(c), emphasizes the reflective characteristics of different objects, rendering it more appropriate for rural roads with a few distinct objects.
@@ III-D2 Methods
Occupancy-based methods [140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 151] utilize information such as the presence or absence of objects within a grid cell to make occupancy predictions. Traditionally, occupancy grid mapping is performed using an Inverse Sensor Model (ISM) and probabilistic methods (e.g., Bayesian filters [152]) to estimate occupancy probability [146, 153]. Additionally, Extended Kalman Filter (EKF) approaches take into account both measurement and motion models, modeling the environment as a Gaussian random field and estimating the state of each cell using a recursive Bayesian filter [154]. Occupancy-based methods provide a concise environmental representation, simplifying obstacle detection and localization for efficient path planning [155]. However, occupancy-based methods may struggle to estimate occupancy states for cells with cluttered or overlapping radar returns [156]. Dynamic objects also pose problems due to their changing nature and intermittent radar measurements.
@@
Amplitude-based methods (e.g., [140, 141, 147, 157]) utilize the measured radar signal reflected intensity information for object detection and classification. Representative algorithms group radar measurements with similar amplitudes to identify potential obstacles. Clustering techniques (e.g., DBSCAN, K-means) are often used in this context [147]. In addition, convolutional neural networks are used as the classifier to distinguish between vehicles and non-vehicles in radar amplitude-based grid maps [141]. Amplitude-based methods offer robustness against clutter and noise, as they exploit variations in radar return amplitudes for object detection. However, amplitude-based representations can be sensitive to sensor noise and uncertainties, as they rely on accurate intensity measurements. The performance of amplitude-based methods may also be influenced by factors like object occlusions, multi-path interference, and the radar’s limited angular resolution.

### v3

@@ I Introduction
As depicted in Fig. 1, radar sensors have extensive applications in autonomous driving, serving various purposes in the perception of the surrounding environment. Radars mounted on the roof, front and angular directions of an intelligent vehicle are utilized to detect objects in the path and surroundings, providing essential information for path planning and adaptive cruise control [7, 8, 9]. The radar inside the cabin is leveraged to monitor occupant vital signs, left-behind children and driver behaviors (e.g., drowsiness, dangerous maneuvers, health emergencies) in a privacy-protected manner, allowing the autonomous driving system to make alerts as well as safety measures [10, 11, 12, 13]. Furthermore, radar systems deployed alongside roadways can measure the speed of passing vehicles and estimate traffic flow. These traffic data are valuable for traffic optimization, congestion management and intelligent transportation systems, leading to improved road safety and smoother traffic operations [14, 15, 16].
@@ II-A Working Principles
Fig. 3 demonstrates a basic radar system that comprises various analog components, including a synthesizer, a Power Amplifier (PA), transmitters (TX antennas), receivers (RX antennas), a Low Noise Amplifier (LNA), a mixer and a Low Pass Filter (LPF), as well as digital components such as an Analog-to-Digital Converter (ADC), a Digital Signal Processor (DSP) and a MicroController Unit (MCU). When the radar starts working, first, the synthesizer generates a linear frequency-modulated pulse called a “chirp”, which is amplified in power and transmitted by the TX antenna. Second, the RX antenna captures the reflected chirp of the target that is amplified with low noise. Third, by combining the RX and TX signals, the mixer produces an Intermediate Frequency (IF) signal, which is then converted into digital values via the ADC. Generally, a radar system contains multiple TX and RX antennas, resulting in multiple IF signals. Information about the target object, such as range, Doppler velocity and azimuth angle, is contained in these IF signals, which can be separated by the DSP using embedding different signal processing algorithms [18]. Based on extracted target information from the DSP and dynamic information from the radar, the MCU serves as a computer to evaluate system requirements and make informed decisions.
@@
In pipeline systems, the number of steps is a critical factor that directly impacts complexity, performance, modularity, and maintainability. A larger number of steps enhances modularity by breaking down complex tasks into smaller, more manageable components, which improves maintainability and facilitates debugging. However, an excessive number of steps can increase computational overhead, introduce latency, and complicate system integration. These trade-offs must be carefully balanced based on the specific requirements of the application.
@@ II-B1 Parameter Estimation
Range ${R}=\frac{cf_{b}}{2S}$ (1) ${R_{res}}=\frac{c}{2B}$ (2) ${c}$ is the speed of the light ($3*10^{8}m/s$). ${f_{b}}$ is the instantaneous frequency difference at the mixers from TX-chirp and RX-chirp. ${S}$ is the slope of a chirp. ${B}$ refers to the bandwidth.
Doppler Velocity ${V}=\frac{\lambda\Delta\phi}{4\pi T_{c}}$ (3) ${V_{res}}=\frac{\lambda}{2T_{f}}$ (4) $\lambda$ is the wavelength of the transmitted signal. ${\Delta\phi}$ is a shift in phase. ${T_{c}}$ is the duration between the chirps. ${T_{f}}$ is cycle time of a frame.
@@ II-B2 RCS Measurement
In terms of mathematical calculations, RCS is a metric that quantifies the ratio between the scattered density in the direction of the radar and the power density intercepted by the object. Since the power is distributed over a sphere, only a small part of this ($4\pi r^{2}$) can be received by the radar. Hence, the expression for RCS takes the form in Equation 9:
@@ III Radar Data Representations
Fig. 6 illustrates the generation process of five different radar data representations. Initially, the raw output from the radar sensor, known as the ADC signal, which is a digital representation of the analog waveforms received by the radar antenna. The ADC signal is a complex and high-dimensional dataset that contains a mixture of target returns, noise, and clutter, which is hardly interpretable by human observers. Thus, researchers utilize 3D FFTs across the sample, chirp, and antenna dimensions to transform the ADC signal into an image-like representation called the radar tensor. The radar tensor organizes the data into a multi-dimensional array where each point corresponds to a specific range, Doppler shift, and angle, making it easier to visualize and analyze the radar returns. Subsequently, peak detection is employed on the radar tensor to eliminate clutter, resulting in a sparse, point-like representation referred to as the point cloud. Each point in the cloud represents a potential object in the environment, with its coordinates corresponding to the object’s range, azimuth, and elevation. By accumulating point clouds over a specific duration and applying grid mapping methods, the grid map representation is generated for the purpose of identifying static objects. Each cell in the grid map contains information about the occupancy or probability of occupancy of that particular location in the environment. Moreover, some researchers perform a Time-Frequency transform following the Range-FFT to extract the micro-Doppler signature, which is utilized for recognizing objects characterized by micro-motion features. The micro-Doppler signature is vital for the recognition of objects with discernible micro-motion patterns, enhancing the capabilities of radar systems in various surveillance and identification scenarios.
@@
V2X-Radar [68] Point Cloud 2024 Object Detection Radar (Oculii Eagle, Arbe Phoenix), Camera, LiDAR Roads (various weather, time and intersection scenarios)
@@ III-A2 Methods
The vital sign-based methods focus on monitoring the physiological signals of occupants inside the vehicle cabin, including heart rate, respiration rate, and presence detection [85, 11, 76, 86, 87]. Signal processing techniques, such as spectrogram analysis, Short Time Fourier Transform (STFT) and Wavelet Transform (WT), are employed to extract the vital sign signals from the radar ADC signal [10, 87]. Machine learning algorithms, such as statistical classifiers or regression models, are trained using labeled data to estimate vital sign parameters from the extracted features [10].
Radar systems used for cabin vital sign monitoring are generally designed for close-range operation, such as within a vehicle or confined space. They focus on high-resolution detection of small movements and physiological changes, and use less power and have smaller form factors than road-oriented radar systems. However, vital sign-based methods using machine learning algorithms have limitations in real-time processing, which is crucial in low-latency applications [88, 89].
@@ III-A3 Discussion
Another limitation is that ADC signals require a developer kit for data acquisition, which makes data processing challenging. Real-time object perception and vital sign monitoring pose computational challenges, requiring efficient hardware and software solutions. Moreover, neural networks processing ADC signals are suitable for classification tasks in closed environments, such as determining the presence or absence of objects and distinguishing between different types of objects [90]. However, in open environments, various distracting factors lead to inaccurate object feature identification.
@@ III-B2 Methods
2D tensor-based methods typically process 2D input radar data by considering range-Doppler and range-azimuth information. To conduct object classification or detection on radar tensors, image-based network architectures (e.g., Faster R-CNN [93], ResNet [94], U-Net [95]) are applied or modified, as has been demonstrated in various studies using RA tensors [91, 96, 4, 97, 98, 99] or RD tensors [100, 101, 102, 99]. For example, Gao et al[91] initiated the object detection process by defining a fixed-size bounding box for each detected object on the RA tensor. In addition, they extracted the radar data within the bounding box from multiple RA tensors and arranged this data in the form of radar data cubes. With the temporal information and object movement pattern, the data cubes serve as features for classifying different objects, such as pedestrians, cars and cyclists. Dong et al[96] applied ResNet [94] architecture on RA tensors for object detection. They also introduced uncertainty estimation for oriented bounding box localization further to enhance the accuracy of the object detection process. In general, 2D tensor-based methods employ a simplified radar data representation, reducing computational complexity. They can exploit correlations between two dimensions to improve detection and localization performance. However, these methods only employ part of the available 3D spatial information from the radar. With limited representation, the accuracy of complex scenarios with multiple objects or occlusions may be reduced.
@@
Compared to using 2D tensor, 3D tensor-based and 4D tensor-based methods offer a more comprehensive representation of radar data, leveraging the complete spatial information and allowing for precise object detection, segmentation, and localization. These methods can effectively handle complex scenarios involving multiple targets, partial occlusions, and varying azimuth angles. However, the increased dimensions of the tensor result in higher computational complexity and system delays.
@@ III-B3 Discussion
Although radar tensors retain more comprehensive information about an object, they also reserve noise and clutter information, which limit the ability to capture subtle object characteristics or distinguish between similar objects. Furthermore, the radar tensor provides a condensed representation of radar data, reducing the overall data size compared to raw ADC signals. Nevertheless, the radar tensor representation still requires considerable memory storage as well as a large bandwidth, especially for the new 4D radar tensors.
@@ III-C1 Datasets
Conventional 3D radar sensors generate sparse point clouds, as observed in datasets such as nuScenes [55], Zender [44], SeeingThroughFog [57], HawkEye [110], AIODrive [58], RADIal [42], RadarScenes [59], aiMotive [63] and MiliPoint [67]. Recently, 4D radar datasets have emerged, including Astyx [111], VoD [61], TJ4DRadSet [62], WaterScenes [64], Dual-Radar [66] and V2X-Radar [68]. While Astyx [111] stands as the first 4D radar point cloud dataset, it is limited by a small data size of only 500 frames. VoD [61] and TJ4DRadSet [62] datasets represent notable advancements in terms of data categories and data size. Meanwhile, these two datasets also incorporate simultaneous LiDAR data, facilitating comparative analysis between 4D radar point clouds and LiDAR point clouds. Interestingly, WaterScenes [64] is a 4D radar dataset focused on objects in waterway environments, which demonstrates the robustness of using 4D radar on water surfaces, particularly in adverse lighting and weather conditions. Dual-Radar [66] incorporates two types of 4D radars, expanding the comparison of different radar performance and further research on practical 4D radar perception algorithms. V2X-Radar [68] dataset integrates the on-vehicle 4D radar with the roadside 4D radar to achieve the Vehicle-to-Everything (V2X) cooperative perception in autonomous driving.
@@ III-C2 Methods
Point-based methods [116, 112, 117, 113, 5, 123, 133] process radar point clouds at the individual point level, extracting spatial features directly from these points. These methods generally draw inspiration from LiDAR-based algorithms, such as PointNet [134], PointNet++ [135] and Frustum PointNets [136]. These techniques leverage neural networks with convolutional filters to process radar points directly, considering spatial coordinates as integral features in the network structure. Schumann et al[116] proposed a structure based on PointNet++ for semantic segmentation on radar point clouds. Specifically, the Multi-Scale Grouping (MSG) module in PointNet++ helps group and generate features for a center point and its neighborhoods. They demonstrated that incorporating the RCS value and compensated Doppler velocity significantly improved classification accuracy. Inspired by [116], Danzer et al[112] adopted a two-stage method using PointNets [134, 136] for 2D car detection and segmentation. They considered each point as a proposal and adjusted the proposal size according to the object’s prior knowledge. Then, PointNet and Frustum PointNets are used to classify the proposals and each point in the proposal. Finally, the bounding box prediction is executed only for the proposals that are objects. Point-based methods are good at accurate localization and classification of objects by considering individual radar points. However, they are susceptible to noise and occlusion, which affect feature extraction accuracy.
@@
Voxel-based methods [137, 114, 138, 139] transform radar point clouds into volumetric representations by discretizing the 3D space into a grid of voxels. Then, 3D convolutions or sparse convolutional neural networks (e.g., VoxelNet [140], SECOND [141]) are leveraged to process the voxels to extract features for perception tasks. For example, GRIF Net [137] initially converts point clouds into voxels. As point clouds typically exhibit sparsity with numerous empty voxels, GRIF Net utilizes the Sparse Block Network (SBNet) [142] to convolve only on masked regions, thereby avoiding unnecessary computations on ineffective blank areas. In LXL [139], the input radar point cloud is initially pillarized as in PointPillars [143]. Then, the pillar representation is fed into SECOND to extract multi-level BEV features with spatial and contextual information. Above all, voxel is an efficient representation for processing dense radar point clouds and is robust to varying point densities and occlusion. However, voxel-based methods lose fine-grained spatial information due to voxelization and discretization. Moreover, the limited resolution of voxel grids lead to decreased accuracy in object detection, especially for small or distant objects.
@@ III-C3 Discussion
Point cloud representation provides an intuitive 3D spatial structure of the surroundings. Since radar point clouds are generated by filtering techniques, they have advantages in filtering noise and being a lightweight data representation. However, some potential information within the raw data inevitably be lost [3]. Additionally, some small objects or objects with weak reflections may not be rendered as point clouds.
@@ III-D Grid Map
Leveraging multi-frame point clouds, some researchers transform 3D radar reflections into a BEV pseudo-image called the grid map [116]. While the grid map is presented as aerial views, it is intended to illustrate the complexity and diversity of environments that autonomous vehicles are navigating. There are two primary types of grid maps commonly used in radar-based perception: one is the occupancy-based grid map [144, 145, 146, 147, 148, 149, 150, 151, 152, 153], which represents the obstacles and free-space information; the other is the amplitude-based grid map [144, 145], which displays the reflected power values in that particular area. As shown in Fig. 9(b), an occupancy-based grid map has a clear outline, making it more feasible for outline-based detection tasks (e.g., parking spaces). An amplitude-based grid map, as indicated in Fig. 9(c), emphasizes the reflective characteristics of different objects, rendering it more appropriate for rural roads with a few distinct objects.
@@ III-D2 Methods
Occupancy-based methods [144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 155] utilize information such as the presence or absence of objects within a grid cell to make occupancy predictions. Traditionally, occupancy grid mapping is performed using an Inverse Sensor Model (ISM) and probabilistic methods (e.g., Bayesian filters [156]) to estimate occupancy probability [150, 157]. Additionally, Extended Kalman Filter (EKF) approaches take into account both measurement and motion models, modeling the environment as a Gaussian random field and estimating the state of each cell using a recursive Bayesian filter [158]. Occupancy-based methods provide a concise environmental representation, simplifying obstacle detection and localization for efficient path planning [159]. However, occupancy-based methods struggle to estimate occupancy states for cells with cluttered or overlapping radar returns [160]. Dynamic objects also pose problems due to their changing nature and intermittent radar measurements.
@@
Amplitude-based methods (e.g., [144, 145, 151, 161]) utilize the measured radar signal reflected intensity information for object detection and classification. Representative algorithms group radar measurements with similar amplitudes to identify potential obstacles. Clustering techniques (e.g., DBSCAN, K-means) are often used in this context [151]. In addition, convolutional neural networks are used as the classifier to distinguish between vehicles and non-vehicles in radar amplitude-based grid maps [145]. Amplitude-based methods offer robustness against clutter and noise, as they exploit variations in radar return amplitudes for object detection. However, amplitude-based representations can be sensitive to sensor noise and uncertainties, as they rely on accurate intensity measurements. The performance of amplitude-based methods is also influenced by factors like object occlusions, multi-path interference, and the radar’s limited angular resolution.